# Two failure modes of deep transformers and how to avoid them: a unified theory of signal propagation at initialisation

### v2

@@ B.2.2 Computation of $Y_{p}(\beta)$
$=\mathbb{E}\left[\frac{\sum\limits_{\sigma}(a_{v\sigma}+a_{u\sigma})e^{h_{\sigma\sigma}(a_{v\sigma}+a_{u\sigma})}}{\sum\limits_{\sigma}e^{h_{\sigma\sigma}(a_{v\sigma}+a_{u\sigma})}}\cdot\frac{\sum\limits_{\sigma}e^{h_{\sigma\sigma}(a_{v\sigma}+a_{u\sigma})}}{\sum\limits_{s,s^{\prime}}e^{h_{ss^{\prime}}(a_{vs}+a_{us^{\prime}})}}\right]$ (B.8)
@@
$Y_{p}(\beta)$ $=\frac{\displaystyle\lim_{\mathbf{h}\to\mathbf{1}}\sum\limits_{\sigma}\partial_{h_{\sigma\sigma}}\mathbb{E}\log\sum\limits_{s,s^{\prime}}e^{h_{ss^{\prime}}(a_{vs}+a_{us^{\prime}})}}{\displaystyle\lim_{h\to 1}\partial_{h}\mathbb{E}\log\sum\limits_{s}e^{h(a_{vs}+a_{us})}}$ (B.9)
@@
$\frac{1}{N}\mathbb{E}\log\sum_{s}e^{h(a_{vs}+a_{us})}=\begin{cases}1+\left(1-\frac{p^{2}}{q^{2}}\right)\beta^{2}h^{2}\quad\beta<\frac{1}{h\sqrt{1-\frac{p^{2}}{q^{2}}}}\\ 2\beta h\sqrt{1-\frac{p^{2}}{q^{2}}}\quad\beta>\frac{1}{h\sqrt{1-\frac{p^{2}}{q^{2}}}}\end{cases}$ (B.10)
@@
$\lim_{h\to 1}\partial_{h}\mathbb{E}\log\sum_{s}e^{h(a_{vs}+a_{us})}=\begin{cases}2\beta^{2}\left(1-\frac{p^{2}}{q^{2}}\right)\quad\beta<\frac{1}{\sqrt{1-\frac{p^{2}}{q^{2}}}}\\ 2\beta\sqrt{1-\frac{p^{2}}{q^{2}}}\quad\beta>\frac{1}{\sqrt{1-\frac{p^{2}}{q^{2}}}}\end{cases}$ (B.11)
@@
$\mathbb{E}\left[\sum_{s,\sigma}A_{ts}A_{t^{\prime}\sigma}q_{s\sigma}\right].$ (B.12)
With the limit $T\to\infty$ in mind, and since we are concerned only with the average overlap, we neglect sub-leading fluctuations and focus directly on the leading terms:
@@
$\mathbb{E}\left[q\sum_{s}A_{ts}A_{t^{\prime}s}+p\sum_{s\neq\sigma}A_{ts}A_{t^{\prime}\sigma}\right].$ (B.13)
@@
$Y^{(2)}(\beta)=\lim_{T\to\infty}\mathbb{E}\left[\sum_{s}A_{ts}^{2}\right],\qquad Y_{p}(\beta)=\lim_{T\to\infty}\mathbb{E}\left[\sum_{s}A_{ts}A_{t^{\prime}s}\right]\quad\text{for }t\neq t^{\prime},$ (B.14)
@@
Substituting the result, we get the update for the average norm:
$q\xleftarrow{\mathcal{S}}p+\left(q-p\right)Y_{q}(\beta)=\begin{cases}p\quad\beta<\beta_{c}(q,p)\\ p+(q-p)\left(1-\frac{\beta_{c}(q,p)}{\beta}\right)\quad\beta>\beta_{c}(q,p)\end{cases}$ (B.15)
@@ B.3 Finite-Size Effects
Here we give a non-rigorous argument on the finite size effects that afflict our asymptotic theory. In the low-$\beta$ regime, the attention is spread approximately uniformly over a number $T^{*}=e^{S(\beta,\rho)}$ of keys, given by an entropic quantity $S(\beta,\rho)=\Phi(\beta,\rho)-\beta\partial_{\beta}\Phi(\beta,\rho)$ (where the free entropy $\Phi$ was defined in Eq.˜B.2). A derivation of the entropy for the REM, very much related to our problem, is explained in depth by [33]. For $\beta<\beta_{c}(\rho)$, this turns out to be:
$S(\beta,\rho)=N\left(1-\frac{\beta^{2}}{\beta_{c}(\rho)}\right)$ (B.16)
@@
$Y^{(2)}(\beta,\rho)\simeq e^{-S(\beta,\rho)}$ (B.17)
@@ B.4 Derivation of Result˜2
$\left\|\frac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right\|_{F}^{2}=\operatorname{tr}\left(\frac{\partial\mathcal{S}(X)}{\partial W_{Q}}\left(\frac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right)^{\top}\right).$ (B.19)
The first part of the derivation parallels the proof of a related result in [10]. However, unlike their approach, we do not assume uniform attention. Instead, we retain the attention explicitly, which allows the previously derived participation ratio to naturally emerge.
@@
$D_{(ij),(kl)}:=\frac{\partial A_{ij}}{\partial a_{kl}}=\delta_{ik}\delta_{jl}A_{ij}-\delta_{ik}A_{ij}A_{il}.$ (B.20)
@@
Now we can substitute the Jacobian components given by eq. (B.20).
@@
$\mathbb{E}\left\|\tfrac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right\|_{F}^{2}=d^{4}\sigma_{K}^{2}\sigma_{V}^{2}\,q\sum_{ij}A_{ij}^{2}(q-p)^{2}+q\sum_{ij}A_{ij}^{3}(q-p)^{2}+q(q-p)^{2}\sum_{ijn}A_{ij}^{2}A_{in}^{2}$ (B.21)
@@
$\frac{1}{d^{2}}\mathbb{E}\left\|\tfrac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right\|_{F}^{2}=\beta\sqrt{\log(T)}\sigma_{v}^{2}\,q(q-p)\Big[(q-p)\big(Y^{(2)}-2Y^{(3)}\big)+p\big(Y^{(2)}\big)^{2}\Big].$ (B.22)
@@
$\frac{1}{d^{2}}\mathbb{E}\left\|\tfrac{\partial\mathcal{L}}{\partial W_{Q}}\right\|_{F}^{2}\leq\mathcal{B}(X)\beta\sqrt{\log(T)}\sigma_{v}^{2}\,q(q-p)\Big[(q-p)\big(Y^{(2)}-2Y^{(3)}\big)+p\big(Y^{(2)}\big)^{2}\Big].$ (B.23)
where $\mathcal{B}(X)$ is a bounded a quantity of $X$ (see the proof of theorem 3.2 [10]).
@@
$\frac{T}{d^{2}\sqrt{\log T}}\mathbb{E}\left\|\tfrac{\partial\mathcal{L}}{\partial W_{Q}}\right\|_{F}^{2}\propto\beta\sigma_{v}^{2}\,q(q-p)\Big[(q-p)\big(Y^{(2)}-2Y^{(3)}\big)+p\big(Y^{(2)}\big)^{2}\Big].$ (B.24)
Since the $Y^{(r)}\to 0$ in the small $\beta$ regime [33], let’s check our predictions for the case where $q=1,\,p\simeq 0$. In this case $\beta_{c}=\sqrt{2}$, but in light of the discussion on finite size effects in Section˜3.2.2, we actually expect our prediction to be sharp up to $\beta_{c}/2\approx 0.7$ and then a crossover between $\beta_{c}/2$ and $\beta_{c}$ to the other solution. This behaviour is correctly predicted, see Fig.˜3(c).
@@ B.5.1 Action of residual connections
B.5.2 Action of MLPs
Finally, we can use the theory developed by [1, 2] to include the effect of the ReLU MLP on signal propagation. For a two-layer ReLU MLP, the propagation of squared norm $q^{(l)}$ and pairwise inner product $p^{(l)}$ across layers is governed by:
$q^{(l)}=\sigma_{w}^{2}\int\mathcal{D}z\,\phi\left(\sqrt{q^{(l-1)}}z\right)^{2}+\sigma_{b}^{2},\quad l=2,\dots,L$ (B.25)
$p^{(l)}=\sigma_{w}^{2}\int\mathcal{D}\mathbf{z}\,\phi\left(\sqrt{q^{(l-1)}}z_{1}\right)\phi\left(\sqrt{q^{(l-1)}}z_{2}\right)+\sigma_{b}^{2}$ (B.26)
where $\phi$ is the ReLU activation function, and $\mathbf{z}=(z_{1},z_{2})^{\top}$ is a pair of standard Gaussian variables with variance 1 and covariance $\rho^{(l-1)}=\nicefrac{{p^{(l-1)}}}{{q^{(l-1)}}}$.
The initial conditions after the first linear layer are $q^{(1)}=\sigma_{w}^{2}+\sigma_{b}^{2}$ and $p^{(1)}=\sigma_{w}^{2}\rho^{(0)}+\sigma_{b}^{2}$. For a two-layer MLP with ReLU activations, the second layer outputs simplify to:
$q^{(2)}=\frac{\sigma_{w}^{2}}{2}q^{(1)}+\sigma_{b}^{2},\qquad p^{(2)}=\frac{\sigma_{w}^{2}}{2}q^{(1)}f(\rho^{(1)})+\sigma_{b}^{2}$ (B.27)
where $f(\rho)$ captures the correlation structure after the ReLU nonlinearity [29],
$f(\rho)=\frac{1}{\pi}\left(\sqrt{1-\rho^{2}}+\rho\left(\pi-\arccos(\rho)\right)\right)$
After adding the final residual connection after the MLP, we find the updated cosine similarity of tokens after a full transformer block:
$\rho^{\text{block}}=\frac{p^{(2)}+\alpha_{\text{MLP}}^{2}\rho^{(0)}}{q^{(2)}+\alpha_{\text{MLP}}^{2}}$ (B.28)
B.5.3 Full Block cosine similarity update algorithms
@@ Appendix C Supplementary Figures
We train gain-controlled transformers with post-LN with skip connections strength $\alpha_{\text{SA}}=\alpha_{\text{MLP}}=1$ of one and twenty layers. The latter case would fail for a standard transformer in the same setting due to rank collapse. The training dynamics is reported in Fig.˜C.1.
Figure C.1: Training 30 layers of vanilla and Gain-controlled Transformer on TinyStories.
@@ C.2 Visualizing the the query/key variance effect on the spectrum of the self-attention matrix
As shown by [34], the spectral bulk of a random stochastic matrix—such as an attention matrix at initialization—has radius $\mathcal{O}(T^{-1/2})$. This result applies under standard initialization schemes, where the variance of the attention scores remains fixed and independent of sequence length.
In contrast, under our proposed rescaling of the query and key weight matrices, the variance of the attention scores scales as $\sigma_{a}^{2}=\log T$, which depends on the dimension of the attention matrix. This places the system in a different spectral regime and, crucially, avoids the spectral bulk to have $\mathcal{O}(T^{-1/2})$ radius, key driver of rank collapse in conventional transformer models, as visualized in Fig.˜C.2.
Figure C.2: Spectrum of a $512\times 512$ self-attention matrix for various values of the query/key variance parameter $\beta$.
@@ C.3 Entropy collapse can be mitigated by low learning rate
Figure C.3: Entropy collapse can be partially mitigated by smaller learning rates.
@@ C.4 Phase transition in infinitely deep transformers
A natural question is whether signal propagation can remain stable at infinite depth. While this is not possible with ReLU activations, it becomes a genuine phenomenon when using tanh. This behaviour was first observed by [1] in MLPs, and more recently extended to Transformers by [13]. In Fig.˜C.4, we confirm the phase transition in forward signal propagation, but emphasise that achieving this behaviour requires placing the MLP in the chaotic phase, which is associated with exploding gradients and unstable training.
Figure C.4: Phase Transition to Infinitely Deep Signal Propagation. (Left) Cosine-similarity update map of a full transformer block with tanh activations in the MLPs. By tuning the MLP variance to enter the chaotic regime, the collapsing effect of self-attention can be counterbalanced, resulting in a non-trivial fixed point in the similarity dynamics. (Right) Iterating the update map reveals the evolution of cosine similarity with depth—predictions align closely with experimental observations.

### v3

@@ B.2.2 Computation of $Y_{p}(\beta)$
$=\mathbb{E}\left[\frac{\sum\limits_{\sigma}(a_{v\sigma}+a_{u\sigma})e^{h_{\sigma\sigma}(a_{v\sigma}+a_{u\sigma})}}{\sum\limits_{\sigma}e^{h_{\sigma\sigma}(a_{v\sigma}+a_{u\sigma})}}\cdot\frac{\sum\limits_{\sigma}e^{h_{\sigma\sigma}(a_{v\sigma}+a_{u\sigma})}}{\sum\limits_{s,s^{\prime}}e^{h_{ss^{\prime}}(a_{vs}+a_{us^{\prime}})}}\right]$ (25)
@@
$Y_{p}(\beta)$ $=\frac{\displaystyle\lim_{\mathbf{h}\to\mathbf{1}}\sum\limits_{\sigma}\partial_{h_{\sigma\sigma}}\mathbb{E}\log\sum\limits_{s,s^{\prime}}e^{h_{ss^{\prime}}(a_{vs}+a_{us^{\prime}})}}{\displaystyle\lim_{h\to 1}\partial_{h}\mathbb{E}\log\sum\limits_{s}e^{h(a_{vs}+a_{us})}}$ (26)
@@
$\frac{1}{N}\mathbb{E}\log\sum_{s}e^{h(a_{vs}+a_{us})}=\begin{cases}1+\left(1-\frac{p^{2}}{q^{2}}\right)\beta^{2}h^{2}\quad\beta<\frac{1}{h\sqrt{1-\frac{p^{2}}{q^{2}}}}\\ 2\beta h\sqrt{1-\frac{p^{2}}{q^{2}}}\quad\beta>\frac{1}{h\sqrt{1-\frac{p^{2}}{q^{2}}}}\end{cases}$ (27)
@@
$\lim_{h\to 1}\partial_{h}\mathbb{E}\log\sum_{s}e^{h(a_{vs}+a_{us})}=\begin{cases}2\beta^{2}\left(1-\frac{p^{2}}{q^{2}}\right)\quad\beta<\frac{1}{\sqrt{1-\frac{p^{2}}{q^{2}}}}\\ 2\beta\sqrt{1-\frac{p^{2}}{q^{2}}}\quad\beta>\frac{1}{\sqrt{1-\frac{p^{2}}{q^{2}}}}\end{cases}$ (28)
@@
$\mathbb{E}\left[\sum_{s,\sigma}A_{ts}A_{t^{\prime}\sigma}q_{s\sigma}\right].$ (29)
With the limit $T\to\infty$ in mind, we neglect sub-leading fluctuations and focus on the leading terms:
@@
This approximation holds due to concentration of measure immediately after the embedding layer. Moreover, the result we derive in the infinite-sequence-length limit for the first layer is independent of the specific token indices, meaning that the simplex geometry is preserved. Consequently, the same reasoning applies recursively at every layer, implying that the dynamics are fully captured by the two scalar quantities $q$ and $p$: they are the only relevant geometric degrees of freedom precisely because the simplex structure persists throughout the network.
$\mathbb{E}\left[q\sum_{s}A_{ts}A_{t^{\prime}s}+p\sum_{s\neq\sigma}A_{ts}A_{t^{\prime}\sigma}\right].$ (30)
@@
$Y^{(2)}(\beta)=\lim_{T\to\infty}\mathbb{E}\left[\sum_{s}A_{ts}^{2}\right],\qquad Y_{p}(\beta)=\lim_{T\to\infty}\mathbb{E}\left[\sum_{s}A_{ts}A_{t^{\prime}s}\right]\quad\text{for }t\neq t^{\prime},$ (31)
@@
Substituting the result, we get the update for the average squared norm:
$q\xleftarrow{\mathcal{S}}p+\left(q-p\right)Y_{q}(\beta)=\begin{cases}p\quad\beta<\beta_{c}(q,p)\\ p+(q-p)\left(1-\frac{\beta_{c}(q,p)}{\beta}\right)\quad\beta>\beta_{c}(q,p)\end{cases}$ (32)
@@ B.3 Finite-Size Effects
Here we give a non-rigorous argument on the finite size effects that afflict our asymptotic theory. In the low-$\beta$ regime, the attention is spread approximately uniformly over a number $T^{*}=e^{S(\beta,\rho)}$ of keys, given by an entropic quantity $S(\beta,\rho)=\Phi(\beta,\rho)-\beta\partial_{\beta}\Phi(\beta,\rho)$ (where the free entropy $\Phi$ was defined in eq.˜19). A derivation of the entropy for the REM, very much related to our problem, is explained in depth by Mézard and Montanari [2009]. For $\beta<\beta_{c}(\rho)$, this turns out to be:
$S(\beta,\rho)=N\left(1-\frac{\beta^{2}}{\beta_{c}(\rho)}\right)$ (33)
@@
$Y^{(2)}(\beta,\rho)\simeq e^{-S(\beta,\rho)}$ (34)
@@ B.4 Derivation of result˜2
$\left\|\frac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right\|_{F}^{2}=\operatorname{tr}\left(\frac{\partial\mathcal{S}(X)}{\partial W_{Q}}\left(\frac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right)^{\top}\right).$ (36)
The first part of the derivation parallels the proof of a related result in Noci et al. [2022]. However, unlike their approach, we do not assume uniform attention. Instead, we retain the attention explicitly, which allows the previously derived participation ratio to naturally emerge.
@@
$D_{(ij),(kl)}:=\frac{\partial A_{ij}}{\partial a_{kl}}=\delta_{ik}\delta_{jl}A_{ij}-\delta_{ik}A_{ij}A_{il}.$ (37)
@@
Now we can substitute the Jacobian components given by eq. (37).
@@
$\mathbb{E}\left\|\tfrac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right\|_{F}^{2}=d^{4}\sigma_{K}^{2}\sigma_{V}^{2}\,q\sum_{ij}A_{ij}^{2}(q-p)^{2}+q\sum_{ij}A_{ij}^{3}(q-p)^{2}+q(q-p)^{2}\sum_{ijn}A_{ij}^{2}A_{in}^{2}$ (38)
@@
$\frac{1}{d^{2}}\mathbb{E}\left\|\tfrac{\partial\mathcal{S}(X)}{\partial W_{Q}}\right\|_{F}^{2}=\beta\sqrt{\log(T)}\sigma_{v}^{2}\,q(q-p)\Big[(q-p)\big(Y^{(2)}-2Y^{(3)}\big)+p\big(Y^{(2)}\big)^{2}\Big].$ (39)
@@
$\frac{1}{d^{2}}\mathbb{E}\left\|\tfrac{\partial\mathcal{L}}{\partial W_{Q}}\right\|_{F}^{2}\leq\mathcal{B}(X)\beta\sqrt{\log(T)}\sigma_{v}^{2}\,q(q-p)\Big[(q-p)\big(Y^{(2)}-2Y^{(3)}\big)+p\big(Y^{(2)}\big)^{2}\Big].$ (40)
where $\mathcal{B}(X)$ is a bounded a quantity of $X$ (see the proof of theorem 3.2 Noci et al. [2022]).
@@
$\frac{T}{d^{2}\sqrt{\log T}}\mathbb{E}\left\|\tfrac{\partial\mathcal{L}}{\partial W_{Q}}\right\|_{F}^{2}\propto\beta\sigma_{v}^{2}\,q(q-p)\Big[(q-p)\big(Y^{(2)}-2Y^{(3)}\big)+p\big(Y^{(2)}\big)^{2}\Big].$ (41)
Since the $Y^{(r)}\to 0$ in the small $\beta$ regime [Mézard and Montanari, 2009], let’s check our predictions for the case where $q=1,\,p\simeq 0$. In this case $\beta_{c}=\sqrt{2}$, but in light of the discussion on finite size effects in section˜3.2, we actually expect our prediction to be sharp up to $\beta_{c}/2\approx 0.7$ and then a crossover between $\beta_{c}/2$ and $\beta_{c}$ to the other solution. This behaviour is correctly predicted, see fig.˜3(c).
@@ B.5.1 Action of residual connections
B.5.2 Full Block cosine similarity update algorithms
@@ Appendix C Supplementary Figures
We train gain-controlled transformers with post-LN with skip connections strength $\alpha_{\text{SA}}=\alpha_{\text{MLP}}=1$ of one and twenty layers. The latter case would fail for a standard transformer in the same setting due to rank collapse. The training dynamics is reported in fig.˜6.
Figure 6: Training 30 layers of vanilla and Gain-controlled Transformer on TinyStories.
@@ C.2 Visualizing the the query/key variance effect on the spectrum of the self-attention matrix
As shown by Bordenave et al. [2012], the spectral bulk of a random stochastic matrix—such as an attention matrix at initialization—has radius $\mathcal{O}(T^{-1/2})$. This result applies under standard initialization schemes, where the variance of the attention scores remains fixed and independent of sequence length.
In contrast, under our proposed rescaling of the query and key weight matrices, the variance of the attention scores scales as $\sigma_{a}^{2}=\log T$, which depends on the dimension of the attention matrix. This places the system in a different spectral regime and, crucially, avoids the spectral bulk to have $\mathcal{O}(T^{-1/2})$ radius, key driver of rank collapse in conventional transformer models, as visualized in fig.˜7.
Figure 7: Spectrum of a $512\times 512$ self-attention matrix for various values of the query/key variance parameter $\beta$.
@@ C.3 Entropy collapse can be mitigated by low learning rate
Figure 8: Entropy collapse can be partially mitigated by smaller learning rates.
@@ C.4 Phase transition in infinitely deep transformers
A natural question is whether signal propagation can remain stable at infinite depth. While this is not possible with ReLU activations, it becomes a genuine phenomenon when using tanh. This behaviour was first observed by Poole et al. [2016] in MLPs, and more recently extended to Transformers by Cowsik et al. [2024]. In fig.˜9, we confirm the phase transition in forward signal propagation, but emphasise that achieving this behaviour requires placing the MLP in the chaotic phase, which is associated with exploding gradients and unstable training.
Figure 9: Phase Transition to Infinitely Deep Signal Propagation. (Left) Cosine-similarity update map of a full transformer block with tanh activations in the MLPs. By tuning the MLP variance to enter the chaotic regime, the collapsing effect of self-attention can be counterbalanced, resulting in a non-trivial fixed point in the similarity dynamics. (Right) Iterating the update map reveals the evolution of cosine similarity with depth—predictions align closely with experimental observations.
@@
C.5 Training Dynamics for the Experiments in the Main Text
Figure 10: Training loss corresponding to the experiment shown in fig.˜1e.
Figure 11: Training loss corresponding to the experiment shown in fig.˜3b.
C.6 Autoregressive Models
All experiments were performed on a $50$-layer autoregressive transformer with post-normalization. We used hidden size $720$, intermediate (MLP) size $720$, and ReLU activations. The weight and bias variances were set to $\sigma^{2}_{w}=\sigma^{2}_{v}=0.2$ and $\sigma^{2}_{b}=4\times 10^{-4}$, respectively. Both the self-attention and MLP blocks employed scaling coefficients $\alpha_{\mathrm{SA}}=1.0$ and $\alpha_{\mathrm{MLP}}=1.0$. Results were averaged over $10$ independent model initializations and $10$ sequences of length approximately $200$ tokens.
Figure 12: Propagation in Autoregressive Models. Cosine similarity between token representations as a function of layer depth for three values of $\beta$. The solid lines show our theoretical predictions, while the markers with error bars show empirical measurements obtained by averaging over all token pairs, across $10$ sequences of length $\approx 200$ and $10$ independent model initialisations. Left: For the first $50$ tokens, the simplex geometry is not yet established, and the empirical evolution departs from the two-parameter theory. Middle: Averaging over the full sequence partially restores the structure but still mixes early and late tokens. Right: For the last $50$ tokens, the geometry is well approximated by our theory: error bars are small, indicating strong concentration of cosine similarity, and the empirical curves follow the theoretical predictions. This illustrates that in causal models the simplex structure—and hence rank-collapse dynamics—emerge progressively and are captured accurately only for sufficiently late tokens.